At the market supply of and demand for some goods are described by formula




Qd=3800-15p




Qs=10p+1300

1
Expert's answer
2021-12-19T18:08:07-0500

Given that:

"Q_d=3800-15p"

"Q_s=10p+1300"

At equilibrium quantity demanded is equal to quantity of good supplied at a given price.

Thus

3800−15p=10p+1300

3800-1300=25p

"p=\frac{2500}{25}=100"

Quantity demanded is equals to:

"Q_d=3800-15p"

substituting the value of p we get

"Q_d=3800-(15\times100)=2300"

Quantity of goods supplied is:

"Q_s=(10\times100)+1300=2300"

This proofs the relation that at equilibrium the quantity supplied is equal to the quantity demanded.